OpenStudy (anonymous):

Show that the slope of the line joining the points (1/a,1),(1/b, b/a) is independent of the value a

jimthompson5910 (jim_thompson5910):

Slope: m = (y2 - y1)/(x2 - x1) m = (b/a - 1)/(1/b - 1/a) m = ??? I'll let you simplify

OpenStudy (anonymous):

@jim_thompson5910 what do they mean independent of the value a?

jimthompson5910 (jim_thompson5910):

if you change the value of 'a', it will not change the slope

OpenStudy (anonymous):

well i got the slope = -b does that suffice because it doesn't include the variable a?

jimthompson5910 (jim_thompson5910):

good, only b has an effect on the slope

jimthompson5910 (jim_thompson5910):

'a' is not in the equation m = -b, so 'a' has no effect

OpenStudy (anonymous):

Thaks. Can you help me with some other problems please?

jimthompson5910 (jim_thompson5910):

yeah just one more

OpenStudy (anonymous):

posted as a new question